# MIND-V: Hierarchical Video Generation for Long-Horizon Robotic Manipulation with RL-based Physical Alignment

### v1

@@ 4.3 Ablation Study
To validate the contributions of our framework’s core components, we benchmark our full model against three ablated variants on long-horizon tasks: (a) one trained only with SFT without GRPO post-training (w/o GRPO); (b) one replacing our affordance localizer [affordance-r1] with a YOLO-World [YOLOWorld] and SAM2 [sam2] pipeline (w/o Affordance); and (c) one disabling the test-time optimization mechanism (w/o Staged Rollouts).
Table 3: Ablation study. All experiments are conducted on long-horizon tasks, evaluating visual quality and functional correctness.
@@
Aesthetic
Quality
 $\uparrow$
Imaging
Quality
 $\uparrow$
PFC
Score
 $\uparrow$
Sub-task Avg.
Rate (%) $\uparrow$
(a) w/o GRPO 0.491 0.675 0.419 60.1
(b) w/o Affordance 0.498 0.680 0.436 45.5
(c) w/o Staged Rollouts 0.482 0.671 0.433 32.7
MIND-V (Full Model) 0.504 0.684 0.445 61.3
As shown in Table 3, the full model consistently outperforms all ablated versions, validating the importance of each component. The removal of GRPO significantly degrades the PFC Score, confirming the efficacy of our RL-based alignment for enhancing physical plausibility. Replacing the affordance module leads to a substantial drop in task success rates, highlighting the criticality of functional grounding for successful manipulation. Disabling Staged Rollouts results in the most pronounced performance degradation, underscoring its crucial function in mitigating error accumulation during long-horizon generation.
@@ 5 Conclusion
This work introduces MIND-V, the first cognition-inspired hierarchical framework for generating long-horizon videos of robotic manipulation. The architecture integrates a Semantic Reasoning Hub (SRH), the Behavioral Semantic Bridge (BSB), and the Motor Video Generator (MVG). By decoupling high-level semantic reasoning from low-level pixel synthesis, our model successfully addresses the critical challenges of long-horizon coherence, semantic grounding, and physical plausibility. The efficacy of our approach is further enhanced by two key innovations: a GRPO post-training phase with a Physical Foresight Coherence (PFC) reward that aligns video generation with physical laws; and Staged Visual Future Rollouts, a test-time strategy that mitigates error accumulation in long-horizon task. Comprehensive experiments demonstrate that MIND-V not only achieves state-of-the-art performance but also establishes a scalable and fully autonomous paradigm for generating high-fidelity embodied AI data.
[…]
@@ 6 Analysis of Computational Cost and Hyperparameters
This section provides a detailed analysis of the computational cost and key hyperparameters of the MIND-V framework. We first examine the framework’s scalability with respect to task length and then present an ablation study on the number of rollout samples ($K$) used in our test-time optimization strategy (Section 3.2).
@@ 6.1 Scalability with Task Length
To validate the efficiency of our framework for long-horizon tasks, we measure the generation time and peak VRAM usage while varying the number of sub-tasks from one to three. The key metrics are summarized in Table 4 and visualized in Figure 7.
Our findings confirm two critical properties of the proposed design. First, the total generation time scales linearly with the number of sub-tasks. The average time per sub-task remains constant at approximately 60 seconds, demonstrating that our framework’s computational time scales predictably with task length (This is a reference value using the Gemini-2.5 API [gemini25]. The inference speed of the SRH is influenced by multiple factors, including VLM API response time and network latency. Reported times represent pure inference duration, excluding network transmission latency.). Second, and crucially, the peak VRAM usage remains constant regardless of the number of sub-tasks. As shown by the consistently sized circles in Figure 7, the peak VRAM remains constant at approximately 70 Gb. This constant memory footprint is a direct benefit of our hierarchical and autoregressive design, where memory is allocated for a single sub-task and subsequently reused. This design makes our approach highly memory-efficient for generating very long-horizon videos.
The data in Table 4 also reveals the internal distribution of resources. The video generation is the most resource-intensive stage, accounting for the majority of the execution time (approx. 65-70%) and VRAM (approx. 86%). The SRH planning stage, in contrast, constitutes a smaller and stable overhead.
Table 4: Analysis of computational cost as a function of the number of sub-tasks. We report total time, average time per sub-task, peak VRAM usage, and the percentage distribution of time and VRAM across the SRH planning and MVG generation stages.
No. of Total Time Avg. Time Peak VRAM Time Dist. (%) VRAM Dist. (%)
Sub-tasks (s) $\downarrow$ per Sub-task (s) $\downarrow$ (Gb) $\downarrow$ Plan (SRH) Gen (MVG) Plan (SRH) Gen (MVG)
@@
6.2 Analysis on the Number of Rollout Samples (K)
The Staged Visual Future Rollouts mechanism (Section 3.2) is governed by a key hyperparameter, $K$, which defines the number of candidate videos generated at each sub-task transition. While a larger $K$ increases the probability of finding a successful trajectory, it also incurs greater computational cost. To analyze this trade-off, we conduct an ablation study by varying $K$ from 1 to 5.
The results, visualized in Figure 8, clearly illustrate the relationship between performance gains and computational cost. As shown in the performance radar chart (Figure 8, left), a significant performance uplift is observed as $K$ increases from 1 to 3, demonstrating the effectiveness of the rollout mechanism in filtering out suboptimal trajectories. For instance, the Task Success Rate jumps from a modest 35.2% at $K=1$ to 61.3% at $K=3$, an absolute increase of 26.1%. However, this trend exhibits sharply diminishing returns, with only marginal gains when increasing $K$ from 3 to 5. In stark contrast, the computational cost, particularly Peak VRAM, scales unfavorably with larger $K$ as shown in the bar chart (Figure 8, right). Peak VRAM consumption, for example, nearly doubles from 70.1 Gb at $K=3$ to 122.0 Gb at $K=5$. This analysis confirms that $K=3$ strikes an optimal balance between functional correctness and computational efficiency. Therefore, we adopt $K=3$ as the default setting for all experiments, as it delivers the best performance-per-cost trade-off.
@@ 6.2 Analysis on the Number of Rollout Samples (K)
Figure 8: Analysis of the trade-off for the number of rollout samples ($K$). (Left) The performance radar chart shows that the overall performance area expands significantly up to K=3 but exhibits diminishing returns thereafter. (Right) The cost chart shows that both time and Peak VRAM increase steadily with K, with memory cost escalating significantly. K=3 (highlighted) is chosen as the optimal balance.
Table 5: Ablation study on the number of rollout samples ($K$). We evaluate the impact of varying $K$ on functional correctness, visual quality, and computational cost. The setting $K=3$ (highlighted) achieves the best balance between performance and efficiency.
@@ 7 Dataset Construction
The Supervised Fine-Tuning (SFT) stage (Section 3.3.1) of our training paradigm requires a large-scale dataset of robotic manipulation videos annotated with our structured Behavioral Semantic Bridge (BSB) representation. To this end, we developed an automated pipeline to generate ground-truth BSB annotations from the raw Bridge V2 dataset [bridge] following the data processing protocol established in [robomaster], as illustrated in Figure 9. This pipeline comprises two primary stages: Object Representation Generation and Trajectory Decomposition.
@@ 7.1 Object Representation Generation
This stage generates the Object Representation (segmentation masks) for the manipulated object ($M_{\text{obj}}$) and the robot arm ($M_{\text{rob}}$) by grounding the natural language instruction in the visual scene. For each sub-task video and its corresponding instruction (e.g., “pick up the red block”), the process is as follows:
@@
    Object Identification: A pre-trained Vision-Language Model (VLM), such as Qwen-VL [Qwen-VL], extracts the noun phrase corresponding to the object of manipulation (e.g., “red block”) from the instruction.
@@
    Robot Arm Segmentation: For the robot arm mask ($M_{\text{rob}}$), we use a pre-defined, comprehensive mask of the manipulator, as its visual features are consistent across all tasks in the dataset.
@@ 7.2 Trajectory Decomposition and Phase Segmentation
The trajectories of the robot arm and the manipulated object are determined by the paths of their respective mask centroids during these phases. The Phase Transition Points ($F_{\text{pre}},F_{\text{interact}},F_{\text{post}}$) are defined by the start and end points of the object’s motion. To ensure data quality, failure cases from the automated pipeline, such as incorrect grounding or trajectory errors, are flagged for manual correction by human annotators.
Figure 11: Qualitative comparison on short-horizon tasks. This figure illustrates performance on two distinct single-step instructions. (Top) For “Put the mushrooms into the metal pot,” the baseline (HunyuanVideo [hunyuanvideo]) exhibits physical implausibility, with the mushroom clipping through the pot’s rim. MIND-V, in contrast, generates a physically plausible interaction. (Bottom) For the more abstract instruction “Clean the floor,” the baseline (Wan-14B [wan22]) fails to take any action, demonstrating a lack of semantic grounding. MIND-V correctly interprets the instruction, grasps the cloth, and performs a wiping motion, showcasing its superior planning and reasoning capabilities.
@@ 8 Additional Visual Results
This section presents a comprehensive qualitative evaluation comparing MIND-V against state-of-the-art baseline models. We analyze performance across three distinct regimes, including multi-stage long-horizon manipulation tasks shown in Figure 10, atomic short-horizon interactions illustrated in Figure 11, and generalization to complex out-of-distribution (OOD) scenarios with diverse action primitives as depicted in Figure 13. These visualizations substantiate our quantitative findings and highlight the practical efficacy of our hierarchical architecture.
@@
Long-horizon Tasks    As illustrated in Figure 10, baseline models exhibit a clear breakdown in long-horizon tasks. They not only violate physical common sense within a single sub-task but, more critically, fail to maintain causal consistency across steps. The baseline’s attempt to interact with an object from a previous sub-task highlights a profound failure in long-range planning and semantic grounding. In contrast, MIND-V maintains robust logical coherence throughout the sequence. This success stems from our hierarchical design: the Semantic Reasoning Hub (SRH) explicitly decomposes complex instructions into structured sub-tasks, while the Staged Visual Future Rollouts mechanism (Section 3.2) actively filters physically implausible transitions.
@@
Short-horizon Tasks    The short-horizon examples in Figure 11 further highlight the nuances of our approach. Even when baselines correctly identify target objects, they may generate physically implausible interactions, such as objects spontaneously appearing or levitating without contact, which our GRPO-based physical alignment (Section 3.3.2) successfully mitigates. Furthermore, MIND-V demonstrates superior reasoning on abstract commands (e.g., “Clean the floor”), successfully inferring the correct sequence of primitive actions like grasping a cloth and wiping. This showcases the SRH’s capability to translate high-level, abstract goals into concrete, executable plans—a key differentiator from monolithic models that often fail on such tasks.
Complex OOD Scenarios and Diverse Interactive Actions.    Figure 13 demonstrates the robust generalization capabilities of MIND-V. Panel (a) illustrates the model’s precision in highly cluttered and visually diverse OOD environments. Whether distinguishing a single piece of bread amidst a dense breakfast spread, or operating within an artistic scene resembling an oil painting, the model accurately isolates the target. Crucially, the background remains strictly static despite high-frequency textures. This stability is attributed to the Behavioral Semantic Bridge (BSB), which injects object-specific masks into the diffusion process to separate the manipulated entity from the complex environment. Furthermore, Panel (b) showcases the ability to execute diverse interactions beyond simple pick-and-place tasks. MIND-V successfully manipulates articulated objects, such as opening a cabinet or closing a microwave, and handles deformable materials like folding a tablecloth. These actions require precise kinematic planning and affordance understanding, which are enabled by the SRH’s integration of VLM reasoning with affordance localization.
@@
Figure 13: Generalization to complex scenarios and diverse manipulation skills. Panel (a) demonstrates the robustness of MIND-V in out-of-distribution (OOD) scenarios. The model accurately isolates and manipulates targets in cluttered environments, such as grabbing bread from a full table or picking a dumpling, as well as in stylistically distinct scenes like picking peaches in an artistic setting. The BSB representation ensures precise control without disturbing the background fidelity. Panel (b) highlights a variety of interactive actions where the model leverages affordance-aware reasoning to execute physics-compliant interactions. This includes manipulating articulated objects like opening a cabinet or closing a microwave, as well as handling deformable materials such as folding a tablecloth.
@@
9 Analysis of GRPO Post-Training
To provide insight into the stability and effectiveness of our GRPO post-training phase (Section 3.3.2), we visualize key training metrics on an epoch-by-epoch basis in Figure 12. For clarity, the plots show the average value of each metric at the end of each epoch. The curves illustrate clear, convergent trends for the policy loss (reward signal), approximate KL divergence, and clip fraction.

### v2

@@ 4.4 Ablation Study
Component Contribution Analysis  To evaluate the contribution of each core component, we compare the full MIND-V model against four ablated variants on long-horizon simulation: (a) w/o GRPO: trained solely with SFT, without GRPO post-training stage; (b) Re Affordance w/ YOLO+SAM2: replacing the affordance localizer [affordance-r1] in BSB with a YOLO-World [YOLOWorld] + SAM2 [sam2] pipeline; (c) w/o Staged Rollouts: disabling the test-time staged optimization mechanism; (d) Re Gemini w/ Qwen3-VL: substituting Gemini 2.5 Pro in SRH with Qwen3-VL [qwen3] to test VLM dependence.
Table 4: Ablation study on long-horizon simulation tasks. Visual quality (Aesthetic and Imaging Quality) and functional correctness (PFC Score and subtask Average Success Rate) are reported. Best results are bolded.
@@
Aesthetic $\uparrow$ Imaging $\uparrow$ PFC Score $\uparrow$ subtask Avg. $\uparrow$
(a) w/o GRPO 0.491 0.675 0.429 0.582
(b) Re Affordance w/ YOLO+SAM2 0.498 0.680 0.445 0.455
(c) w/o Staged Rollouts 0.482 0.671 0.438 0.327
(d) Re Gemini w/ Qwen3-VL 0.500 0.679 0.452 0.567
MIND-V (Full) 0.504 0.684 0.462 0.613
As shown in Table 4, the full model outperforms all ablated variants, confirming the necessity of each component. Removing GRPO results in a clear decline in PFC Score, underscoring the effectiveness of our RL-based alignment in improving physical plausibility. Replacing the affordance module with an alternative pipeline causes a substantial drop in subtask success rates, highlighting its critical role in enabling functional grounding for manipulation in BSB. Disabling staged rollouts leads to pronounced performance degradation, demonstrating proposing and verifying futures (“System 2 thinking” [system2]) is essential for mitigating error accumulation over extended horizons. Switching the VLM in SRH yields a marginal drop, indicating that a powerful VLM is beneficial for improving performance, but MIND-V does not rely on a specific VLM model.
Table 5: Sensitivity analysis of reward weights.
$w_{a}$ $w_{p}$ Aesthetic $\uparrow$ Imaging $\uparrow$ PFC Score $\uparrow$ subtask Avg. $\uparrow$
0.5 0.2 0.498 0.651 0.449 0.612
1.0 0.2 0.504 0.658 0.445 0.613
2.0 0.2 0.508 0.657 0.434 0.562
Reward Sensitivity Analysis  These parameters $w_{a}$ and $w_{p}$ in Eq. 3 serve to normalize reward scales to the same order of magnitude. The sensitivity analysis (Table 5) on long-horizon tasks shows robust performance across a reasonable range and the rationality of our parameter selection.
@@ 5 Conclusion
In this work, we introduce MIND-V, a cognitive hierarchical video world model designed to simulate the physical reality of long-horizon robotic manipulation. Drawing inspiration from the human brain’s cognition-to-execution pathway, MIND-V integrates three core modules: the Semantic Reasoning Hub (SRH), the Behavioral Semantic Bridge (BSB), and the Motor Video Generator (MVG). By decoupling high-level semantic reasoning from low-level pixel synthesis, MIND-V successfully addresses the key challenges of long-horizon coherence, semantic grounding, and physical plausibility. The efficacy of our approach is further enhanced by two key innovations: a GRPO post-training phase with a Physical Foresight Coherence (PFC) reward that enforces alignment with physical laws; and a test-time optimization via staged visual future rollouts that mitigates error accumulation over extended horizons. Comprehensive experiments demonstrate that MIND-V not only achieves SOTA performance in video synthesis but also serves as an effective training ground for embodied AI policies.
@@ 6 Analysis of Computational Cost and Hyperparameters
This section provides a detailed analysis of the computational cost and key hyperparameters of the MIND-V framework. We first examine the world model’s scalability with respect to task length and then present an ablation study on the number of staged visual future rollouts (K) utilized in our test-time optimization strategy.
@@ 6.1 Scalability with Task Length
To validate MIND-V’s efficiency as a world simulator in long-horizon tasks, we measure the simulation time and peak VRAM usage across tasks ranging from one to three sub-tasks. The key metrics are summarized in Table 6 and visualized in Figure 7.
Our findings highlight two critical properties of the proposed architecture. First, the total simulation time scales linearly with the number of sub-tasks. The average time per sub-task remains constant at approximately 60 seconds  [^1] This is a reference value using the Gemini-2.5 API [gemini25]. The inference speed of the SRH is influenced by multiple factors, including VLM API response time and network latency. Reported times represent pure inference duration, excluding network transmission latency., demonstrating that our framework’s computational time scales predictably with task length. Second, and crucially, the peak VRAM usage remains constant regardless of the task length. As shown by the consistently sized circles in Figure 7, the peak VRAM remains constant at approximately 70 GB. This constant memory footprint is a direct benefit of our hierarchical and autoregressive design, where memory is allocated for a single sub-task and subsequently reused, enabling highly scalable visual world simulation for long task sequences.
Table 6: Computational cost as a function of task length. We report total time, average time per sub-task, peak VRAM usage, and the percentage distribution of time and VRAM across the SRH planning and MVG simulation stages.
No. of Total Avg. Time per Peak VRAM Time Dist. (%) VRAM Dist. (%)
Sub-tasks Time (s) Sub-task (s) (GB) Plan (SRH) Gen (MVG) Plan (SRH) Gen (MVG)
@@
Furthermore, Table 6 details the internal resource distribution. The Motor Video Generator (MVG) constitutes the primary computational bottleneck, consuming the majority of both execution time (approximately 65-70%) and VRAM (approximately 86%). In contrast, the Semantic Reasoning Hub (SRH) planning stage introduces only a minor and stable computational overhead.
6.2 Analysis of the Number of Rollout Samples (K)
The Staged Visual Future Rollouts mechanism is governed by a key hyperparameter, K, which defines the number of candidate future trajectories simulated at each sub-task transition. While a larger K increases the probability of identifying a physically plausible future, it also incurs greater computational cost. To analyze this trade-off, we conduct an ablation study by varying K from 1 to 5.
@@ 6.2 Analysis of the Number of Rollout Samples (K)
The results, visualized in Figure 8, illustrate the relationship between simulation fidelity and computational cost. As shown in the performance radar chart (Figure 8, left), a significant performance uplift is observed as K increases from 1 to 3, demonstrating the effectiveness of the rollout mechanism in filtering out physically incoherent futures. For instance, the Task Success Rate increases from 35.2% at K=1 to 61.3% at K=3. However, this trend exhibits sharply diminishing returns, with only marginal gains when increasing K from 3 to 5. In contrast, the computational cost scales unfavorably with larger K as shown in the bar chart (Figure 8, right). For example, Peak VRAM consumption nearly doubles from 70.1 GB at K=3 to 122.0 GB at K=5. This analysis confirms that K=3 strikes an optimal balance between simulation accuracy and computational efficiency. Therefore, we adopt K=3 as the default setting for all experiments.
Figure 8: Analysis of the trade-off for the number of visual future rollouts (K). (Left) The performance radar chart shows that the overall performance area expands significantly up to K=3 but exhibits diminishing returns thereafter. (Right) The cost chart shows that both time and Peak VRAM increase steadily with K, with memory cost escalating significantly. K=3 (highlighted) is chosen as the optimal balance.
Table 7: Ablation study on the number of visual future rollouts (K). We evaluate the impact of varying K on functional correctness, visual quality, and computational cost. The setting K=3 (highlighted row) achieves the best balance between performance and efficiency.
@@ 7 Dataset Construction
The Supervised Fine-Tuning (SFT) stage of our training paradigm requires a large-scale dataset of robotic manipulation videos annotated with our structured Behavioral Semantic Bridge (BSB) representation. To this end, we developed an automated pipeline to generate ground-truth BSB annotations from the raw Bridge V2 dataset [bridge] following the data processing protocol established in [robomaster], as illustrated in Figure 9. This pipeline comprises two primary stages: Object Representation Generation and Trajectory Decomposition.
@@ 7.1 Object Representation Generation
This stage derives the Object Representation (segmentation masks) for both the manipulated object ($M_{\text{obj}}$) and the robot arm ($M_{\text{rob}}$) by visually grounding the natural language instructions. For each sub-task video and its corresponding instruction (e.g., “pick up the red block”), the extraction process proceeds as follows:
@@
    Object Identification: A pre-trained Vision-Language Model (VLM), such as Qwen-VL 2.5 [Qwen-VL], extracts the noun phrase corresponding to the object of manipulation (e.g., “red block”) from the instruction.
@@
    Robot Arm Mask: Because the manipulator’s visual appearance remains consistent across the dataset, we apply a pre-defined template mask to obtain the robot arm segmentation ($M_{\text{rob}}$).
@@ 7.2 Trajectory Decomposition and Phase Segmentation
The spatiotemporal trajectories of the robot arm and the manipulated object are determined by the paths of their respective mask centroids during these phases. The Phase Transition Points ($F_{\text{pre}},F_{\text{interact}},F_{\text{post}}$) are defined by the start and end points of the object’s motion. To ensure fidelity of world dynamics, failure cases from the automated pipeline, such as incorrect grounding or trajectory errors, are flagged for manual correction by human annotators.
Figure 11: Qualitative comparison on short-horizon tasks. This figure illustrates performance on two distinct single-step instructions. (Top) For “Put the mushrooms into the metal pot”, the baseline (HunyuanVideo [hunyuanvideo]) exhibits physical implausibility, with the mushroom clipping through the pot’s rim. MIND-V, in contrast, simulates a physically plausible interaction. (Bottom) For the more abstract instruction “Clean the floor”, the baseline (Wan-14B [wan22]) fails to take any action, demonstrating a lack of semantic grounding. MIND-V correctly interprets the instruction, grasps the cloth, and performs a wiping motion, showcasing its superior planning and reasoning capabilities.
@@ 8 Additional Visual Results
This section presents a comprehensive qualitative evaluation comparing MIND-V against state-of-the-art baselines. We analyze performance across three distinct regimes, including multi-stage long-horizon manipulation tasks shown in Figure 10, atomic short-horizon interactions illustrated in Figure 11, and generalization to complex out-of-distribution (OOD) scenarios with diverse action primitives as depicted in Figure 12. These visualizations substantiate our quantitative findings, showcasing MIND-V’s superiority as a physically grounded and logically coherent world simulator.
@@
Long-horizon Tasks.  As illustrated in Figure 10, baseline models exhibit a clear breakdown in long-horizon tasks. They not only violate physical common sense within a individual sub-task but also fail to maintain causal consistency across the temporal sequence. For instance, baseline attempts to interact with an object that spontaneously disappeared in a previous step highlight a profound failure in tracking latent world states. In contrast, MIND-V maintains robust logical coherence throughout the sequence. This success stems from our hierarchical design: the Semantic Reasoning Hub (SRH) explicitly decomposes complex instructions into structured sub-tasks, while the Staged Visual Future Rollouts mechanism actively filters physically implausible transitions, effectively preventing long-term semantic drift.
Figure 12: Generalization to complex scenarios and diverse manipulation skills. Panel (a) demonstrates the robust generalization ability of MIND-V in out-of-distribution (OOD) scenarios. The world model accurately isolates and manipulates targets in cluttered environments (e.g., grabbing bread from a full table or picking a dumpling), as well as in stylistically distinct scenes (e.g., picking peaches in an artistic setting). The domain-invariant BSB representation ensures precise control without disturbing the background fidelity. Panel (b) highlights the simulation of diverse interactive dynamics where the model leverages affordance-aware reasoning to execute physics-compliant interactions. This includes manipulating articulated objects like opening a cabinet or closing a microwave, as well as handling deformable materials such as folding a tablecloth.
@@
Short-horizon Tasks.  The short-horizon examples in Fig. 11 further emphasize the physical fidelity of our approach. Even when baselines correctly identify target objects, they frequently hallucinate physically impossible interactions, such as objects levitating without contact or penetrating solid surfaces. Our GRPO-based physical alignment successfully mitigates these dynamic violations. Furthermore, MIND-V excels at interpreting abstract commands (e.g., “clean the floor”) by autonomously inferring the correct latent action sequence (grasping a cloth, then wiping). This demonstrates the SRH’s cognitive capacity to translate high-level intent into executable, physically grounded trajectories—a key differentiator from monolithic models that struggle with action grounding.
Complex OOD Scenarios and Diverse Interactive Actions.  Fig. 12 demonstrates the robust generalization capabilities that are the hallmark of a scalable world model. Panel (a) illustrates the model’s precision in visually diverse OOD environments. Whether distinguishing a specific object within severe clutter or operating in a stylized scene, MIND-V accurately executes the intent. Crucially, the background remains strictly static despite high-frequency textures. This stability is achieved because the Behavioral Semantic Bridge (BSB) decouples semantic intent from visual appearance, providing structural, domain-invariant guidance to the diffusion process. Panel (b) highlights MIND-V’s ability to model complex physical state transitions. The model effortlessly simulates interactions with articulated objects (e.g., cabinets) and deformable materials (e.g., cloth). Executing these actions requires deep kinematic and affordance understanding, proving that integrating the SRH’s cognitive reasoning with the MVG’s simulation engine creates a highly capable embodied world model.
9 Analysis of GRPO Post-Training
@@
To provide insight into the stability and effectiveness of our GRPO post-training phase, we visualize key training metrics on an epoch-by-epoch basis in Figure 13. For clarity, the plots show the average value of each metric at the end of each epoch. The curves illustrate clear, convergent trends for the policy loss (reward signal), approximate KL divergence, and clip fraction.